OpenStudy (anonymous):

What is the solution to the system of equations {2x+3y=11 {3x+3y=18 Choices 7,-3 7,25/3 7,-1 -7,25/3

OpenStudy (anonymous):

just 2x +3y - (3x + 3y) = 11 - (18) -x = -7 x = 7 2*7 + 3y =11 3y = 11 -14 y = -3/3 y = -1

OpenStudy (anonymous):

thanks for your help I needed it